OpenStudy (anonymous):

True or false Taneeka borrowed 12000 for a car for 6 years at an arp of 7.25% her monthly payment will be 206.03

OpenStudy (anonymous):

i did the formula but i got 72.00

OpenStudy (mertsj):

What formula did you do?

OpenStudy (mathmate):

The answer is correct if the annual interest rate is compounded every month at 7.25/12% per month. I got exactly $206.03 per month.

OpenStudy (anonymous):

monthly payment formula. m=p(r/12)(1+r/12)/(1+r/12)^12t-1

OpenStudy (mertsj):

You are correct.

OpenStudy (anonymous):

@mathmate are u sure? why do i keep getting 72.00 if i work it out?

OpenStudy (mathmate):

I am looking at your formula to spot any errors. If you want, I'll show you mine.

OpenStudy (anonymous):

can u please because i want to learn how to do it right ill show u what i did to so i can see where i went wrong.

OpenStudy (anonymous):

m=12000(.0725/12)(1+.0725/12)^12(6) divded by 1+(.0725/12)^12(6) then i got m=12000(.0060)(1.0060)^72 divided by 1.5429 then i got m=12000(.0060)(1.5385) divded by 1.5429 then finally 110.7611/1.5429=72.00

OpenStudy (mathmate):

My formula is quite simple, Monthly amount =PR^n(R-1)/(R^n-1) P=principal, 12000 R=rate (yearly=1.0725, monthly=1.006041667) n=number of periods (yearly=6, monthly 72) If I calculate yearly compounding where the money is paid monthly, but only credited yearly, it comes to $211.41765 If I calculate monthly credit and compounding, i.e. n=72, R=1+0.0725/12=1.0041667), I get $206.3167

OpenStudy (mathmate):

I looks like your formula is similar to mine, but missing the R^n term.

OpenStudy (mathmate):

correction: R=rate (yearly=1.0725, monthly=1.006041667) should have read: R=rate (yearly=1.0725, monthly=1.0041667)

OpenStudy (mathmate):

Correction: final value for compounding monthly is $206.031673